On the secind mathematicso competition,Delia answered 3÷5 of the problem correctly andzakhensa mother answered 2÷3 of themy correctly. If each problem worth the same mark, who performed better between the two? Show all your working
1
Expert's answer
2019-09-04T14:14:47-0400

Solution:


"\frac{3}{5}=\frac{3\cdot{3}}{5\cdot{3}}=\frac{9}{15}"


"\frac{2}{3}=\frac{2\cdot{5}}{3\cdot{5}}=\frac{10}{15}"


"10>9"


"\frac{10}{15}>\frac{9}{15}"


"\frac{2}{3}>\frac{3}{5}"


Answer:

"\frac{2}{3}>\frac{3}{5}"

Zakhensa performed better then Delia.